# Numerical optimization for the compatibility constant of the lasso

### v1

@@ 6 Concluding remarks
We have provided an approach to compute the compatibility constant when the true zero/nonzero pattern of the true regression coefficients is known. The computation of the compatibility constant reduces to a QP if the signs of the true nonzero regression coefficients are given. Thus, we can compute the compatibility constant by solving QP for all combinations of signs. When the number of nonzero coefficients $s$ is large, we may alternatively use the MIQP with Big-$M$ or SOS1 under a time limit to obtain an upper bound of the compatibility constant. A more efficient computation for large $s$ would be crucial, but we will leave this as a future research topic.
The compatibility condition is one of the most important conditions in the lasso literature for achieving good accuracy in high-dimensional settings. More generally, researchers are interested in the theoretical properties of statistical methods under certain conditions. These conditions must be realistic, but in many cases they cannot be evaluated using a given dataset. This study attempted to address this issue by evaluating a key regularity condition of the lasso, which may serve as a guideline for assessing whether the method is likely to perform well. We believe that methods for checking such conditions using a given dataset would be useful in many practical situations. In the future, we would like to develop numerical approaches to check the regularity conditions and assess whether a method can be expected to perform well.
@@ Appendix A Compatibility constant under compound symmetry
The compatibility constant is closely related to the smallest eigenvalue of the covariance matrix. The eigenstructure of $\Sigma_{\rho}$ is simple: there exists a unique maximum eigenvalue $\lambda_{\max}=1+(p-1)\rho$ with the eigenvector being $\mathbf{1}$, and for the other (p1)(p-1)-dimensional vectors that is orthogonal to $\bm{1}$, we have the same eigenvalues $\lambda_{\min}=1-\rho$. Thus, the quadratic form $\bm{v}^{\top}\Sigma_{\rho}\bm{v}$ is smallest when $\bm{v}$ lies close to the subspace orthogonal to $\mathbf{1}$, that is, when $\sum_{j=1}^{p}v_{j}\approx 0$. This is also reasonable from Eq. (15) as the second term penalizes the component along $\mathbf{1}$.
@@
$\begin{aligned} \phi^{2}&=1-\rho&\quad&(s:\ \text{even}),\\ \phi^{2}&\leq(1-\rho)\left(1+\frac{1}{sr}\right)&\quad&(s:\ \text{odd}),\end{aligned}$ (17)
@@
Next, we derive an explicit upper bound of $\phi$ when $s$ is odd. It is difficult to define $\bm{v}$ as in (18) becuase $s$ is odd, but we may define $\bm{v}$ that is similar to that in the even case in (18):

### v2

@@ 6 Concluding remarks
We have provided an approach to compute the compatibility constant when the support of the true regression coefficients is known. The computation of the compatibility constant reduces to a QP if the signs of the true nonzero regression coefficients are given. Thus, we can compute the compatibility constant by solving QP for all combinations of signs. When the number of nonzero coefficients $s$ is large, we may alternatively use the MIQP with Big-$M$ or SOS1 under a time limit to obtain an upper bound of the compatibility constant. A more efficient computation for large $s$ would be crucial, but we will leave this as a future research topic.
The compatibility condition is one of the most important conditions in the lasso literature for achieving good prediction accuracy in high-dimensional settings. Although the theoretical upper bound is useful, it can be conservative in finite samples. Therefore, if we are interested in the finite-sample behavior of the lasso, both the condition and its implications should be interpreted carefully. More generally, researchers are interested in developing strong theoretical properties of statistical methods under conditions that are as mild as possible. In many cases, however, these conditions cannot be evaluated with a given dataset due to the fact that the true parameter is unknown and the associated optimization problem is often nonconvex. Although exact verification is impossible in practice, it is still meaningful to consider numerical approaches to assess whether the regularity conditions are likely to hold for a given dataset. The evaluation of regularity conditions would be useful in many practical situations, as it helps determine whether the method is expected to perform well. In the future, we would like to develop numerical approaches to evaluate the regularity conditions for various statistical methods.
@@ Appendix A Compatibility constant under compound symmetry
The compatibility constant is closely related to the smallest eigenvalue of the covariance matrix. The eigenstructure of $\Sigma_{\rho}$ is simple: there exists a unique maximum eigenvalue $\lambda_{\max}=1+(p-1)\rho$ with the eigenvector being $\mathbf{1}_{p}$, and for the other (p1)(p-1)-dimensional vectors that is orthogonal to $\bm{1}$, we have the same eigenvalues $\lambda_{\min}=1-\rho$. Thus, the quadratic form $\bm{v}^{\top}\Sigma_{\rho}\bm{v}$ is smallest when $\bm{v}$ lies close to the subspace orthogonal to $\mathbf{1}_{p}$, that is, when $\sum_{j=1}^{p}v_{j}\approx 0$. This is also reasonable from Eq. (13) as the second term penalizes the component along $\mathbf{1}_{p}$.
@@
$\begin{aligned} \phi^{2}&=1-\rho&\quad&(s:\ \text{even}),\\ \phi^{2}&\leq(1-\rho)\left(1+\frac{1}{sr}\right)&\quad&(s:\ \text{odd}),\end{aligned}$
@@
Next, we derive an explicit upper bound of $\phi$ when $s$ is odd. It is difficult to define $\bm{v}$ as in (15) because $s$ is odd, but we may define $\bm{v}$ that is similar to that in the even case in (15):
@@
Appendix B Derivation of the MSPE bound in Proposition 4
Here, we derive the upper bound in Proposition 4 step by step and clarify which inequality is used at each stage. Our goal is to derive
$\frac{1}{n}\|X\hat{\bm{\beta}}-X\bm{\beta}\|_{2}^{2}\leq\frac{9\lambda^{2}s}{\phi^{2}},$
where $\phi$ denotes the compatibility constant.
Step 1: Basic inequality.
Since $\hat{\bm{\beta}}$ minimizes the lasso objective function, comparing the objective values at $\hat{\bm{\beta}}$ and $\bm{\beta}$ yields
$\frac{1}{2n}\|\bm{y}-X\hat{\bm{\beta}}\|_{2}^{2}+\lambda\|\hat{\bm{\beta}}\|_{1}\leq\frac{1}{2n}\|\bm{y}-X\bm{\beta}\|_{2}^{2}+\lambda\|\bm{\beta}\|_{1}.$ (17)
Eq. (17) is referred to as the basic inequality. Let $\bm{\Delta}$ be
$\bm{\Delta}=\hat{\bm{\beta}}-\bm{\beta}.$ (18)
Rearranging Eq. (17) gives
$\frac{1}{2n}\|X\bm{\Delta}\|_{2}^{2}\leq\frac{1}{n}\bm{\varepsilon}^{\top}X\bm{\Delta}+\lambda\bigl(\|\bm{\beta}\|_{1}-\|\hat{\bm{\beta}}\|_{1}\bigr).$ (19)
Step 2: Dual norm
For given vectors $\bm{a}$ and $\bm{b}$ with the same dimension, we have $|\bm{a}^{T}\bm{b}|\leq\|\bm{a}\|_{\infty}\|\bm{b}\|_{1}$. Therefore, we obtain
$\frac{1}{n}\bm{\varepsilon}^{\top}X\bm{\Delta}=\left(\frac{X^{\top}\bm{\varepsilon}}{n}\right)^{\top}\bm{\Delta}\leq\left\|\frac{X^{\top}\bm{\varepsilon}}{n}\right\|_{\infty}\|\bm{\Delta}\|_{1}.$ (20)
Eq. (20) corresponds to the step where the stochastic term is bounded by the maximum absolute correlation.
Step 3: Max bound
Define the event
$\mathcal{T}=\left\{\left\|\frac{X^{\top}\bm{\varepsilon}}{n}\right\|_{\infty}\leq\frac{\lambda}{2}\right\}.$ (21)
Suppose that $\lambda$ satisfies $\lambda\geq 2\sigma\sqrt{2/n(1+\log\left(\frac{p}{\delta}\right))}$ with $\delta\in(0,1)$. Then we have $P(\mathcal{T})\geq 1-\delta$ (e.g., Bühlmann and van de Geer, 2011). When $\mathcal{T}$ holds, Eq. (19) implies
$\frac{1}{2n}\|X\bm{\Delta}\|_{2}^{2}+\lambda\|\hat{\bm{\beta}}\|_{1}\leq\frac{\lambda}{2}\|\bm{\Delta}\|_{1}+\lambda\|\bm{\beta}\|_{1}.$ (22)
Step 4: Triangle inequality
Since $\bm{\beta}_{S^{c}}=0$, Eq. (18) implies
$\bm{\Delta}_{S^{c}}=\hat{\bm{\beta}}_{S^{c}},\qquad\|\bm{\Delta}\|_{1}=\|\bm{\Delta}_{S}\|_{1}+\|\bm{\Delta}_{S^{c}}\|_{1}.$ (23)
By the triangle inequality,
$\|\hat{\bm{\beta}}_{S}\|_{1}=\|\bm{\beta}_{S}+(\hat{\bm{\beta}}_{S}-\bm{\beta}_{S})\|_{1}\geq\|\bm{\beta}_{S}\|_{1}-\|\hat{\bm{\beta}}_{S}-\bm{\beta}_{S}\|_{1}=\|\bm{\beta}_{S}\|_{1}-\|\bm{\Delta}_{S}\|_{1}.$ (24)
Thus, we obtain
$\|\hat{\bm{\beta}}\|_{1}\geq\|\bm{\beta}_{S}\|_{1}-\|\bm{\Delta}_{S}\|_{1}+\|\hat{\bm{\beta}}_{S^{c}}\|_{1}.$ (25)
Substituting Eqs. (23) and (25) into Eq. (22), we obtain
$\frac{1}{n}\|X\bm{\Delta}\|_{2}^{2}\leq 3\lambda\|\bm{\Delta}_{S}\|_{1}-\lambda\|\bm{\Delta}_{S^{c}}\|_{1}.$ (26)
Step 5: Drop negative term
By dropping the nonnegative term $\lambda\|\bm{\Delta}_{S^{c}}\|_{1}$ from the right-hand side of Eq. (26), we obtain
$\frac{1}{n}\|X\bm{\Delta}\|_{2}^{2}\leq 3\lambda\|\bm{\Delta}_{S}\|_{1}.$ (27)
Step 6: Compatibility condition
Eq. (26) implies that $\bm{\Delta}$ satisfies
$\|\bm{\Delta}_{S^{c}}\|_{1}\leq 3\|\bm{\Delta}_{S}\|_{1}.$
Applying the compatibility condition in (2) gives
$\|\bm{\Delta}_{S}\|_{1}^{2}\leq\frac{s\,\bm{\Delta}^{\top}\hat{\Sigma}\bm{\Delta}}{\phi^{2}}=\frac{s}{\phi^{2}}\cdot\frac{1}{n}\|X\bm{\Delta}\|_{2}^{2}.$
Taking square roots on both sides and using Eq. (27), we obtain
$\frac{1}{n}\|X\bm{\Delta}\|_{2}^{2}\leq 3\lambda\sqrt{\frac{1}{n}\|X\bm{\Delta}\|_{2}^{2}\cdot\frac{s}{\phi^{2}}}.$ (28)
Step 7: Final step
Squaring both sides of (28) and rearranging gives the desired upper bound:
$\frac{1}{n}\|X\bm{\Delta}\|_{2}^{2}\leq\frac{9\lambda^{2}s}{\phi^{2}}.$